OpenStudy (anonymous):

find the length of c if the measure of angle A=62 degrees, the measure of angle B=39 degrees, b=15

OpenStudy (anonymous):



OpenStudy (anonymous):

find C first easy since the interior angles have to add up to 180 then use the law of sines

OpenStudy (anonymous):

\[\frac{c}{\sin(82)}=\frac{15}{\sin(39)}\iff c=\frac{15\sin(82)}{\sin(39)} \]

OpenStudy (anonymous):

ok so C=81

OpenStudy (anonymous):

right so i had a typo above

OpenStudy (anonymous):

should be \[\frac{c}{\sin(81)}=\frac{15}{\sin(39)}\iff c=\frac{15\sin(81)}{\sin(39)}\]